# Isometric embeddings of surfaces for scl

### v2

@@ 5. Isometries for scl and the relative Gromov seminorm
Hence, $\Sigma$ has a disc component consisting of $H$ and its two incident vertex discs (note that these two discs are distinct by connectedness of their link). This is impossible since admissible surfaces in standard form are assumed to be disc-free. ∎
@@
In Theorem A, observe that $H_{1}$-injectivity of $T$ could be replaced by the equivalent assumption that $H_{2}(S,T)=0$. With that assumption, the theorem would also hold when $S$ is closed since in that case, $H_{2}(S,T)=0$ implies that $S=T$ by Corollary 3.6. However, this would only add a trivial statement and therefore the generality of the theorem would not be increased.
@@ 6. Extremal surfaces and quasimorphisms
Note however that, even if extremal surfaces were known to exist for the relative Gromov seminorm, we would not obtain an analogue of Corollary 6.1 in that setting. Indeed, to prove Theorem B, we needed to apply Theorem 5.1 with assumption (ii) and work with admissible surfaces in perfect standard form. But an asymptotic promotion argument was necessary to obtain the perfect standard form (see §4.d), and this does not preserve extremal surfaces.
@@
It is natural at this point to ask whether it is equivalent for a chain $c\in B_{1}(\pi_{1}T;\mathbb{Z})$ to rationally bound a positive immersed surface in $T$ or in $S$. Using the same kind of argument as in the proof of Proposition 6.5, it is easy to see that this is true if the word ‘immersed’ is replaced by ‘embedded’. In general, using Scott’s theorem on subgroup separability of surface groups [scott1, scott2], we can lift an immersed surface to an embedded one, and thus obtain an affirmative answer:

### v3

@@ 5. Isometries for scl and the relative Gromov seminorm
Hence, $\Sigma$ has a disc component consisting of $H$ and its two incident vertex discs (note that these two discs are distinct by connectedness of their link). This is impossible since admissible surfaces in standard form are assumed to be disc- and sphere-free. ∎
@@
In Theorem A, observe that $H_{1}$-injectivity of $T$ could be replaced with the equivalent assumption that $H_{2}(S,T)=0$. With that assumption, the theorem would also hold when $S$ is closed since in that case, $H_{2}(S,T)=0$ implies that $S=T$ by Corollary 3.6. However, this would only add a trivial statement and therefore the generality of the theorem would not be increased.
@@ 6. Extremal surfaces and quasimorphisms
Note however that, even if extremal surfaces were known to exist for the relative Gromov seminorm, we would not obtain an analogue of Corollary 6.1 in that setting. Indeed, to prove Theorem B, we needed to apply Theorem 5.1 with assumption (ii) and work with admissible surfaces in perfect standard form. But an asymptotic promotion argument was necessary to obtain the perfect standard form (see §4.e), and this does not preserve extremal surfaces.
@@
It is natural at this point to ask whether it is equivalent for a chain $c\in B_{1}(\pi_{1}T;\mathbb{Z})$ to rationally bound a positive immersed surface in $T$ or in $S$. Using the same kind of argument as in the proof of Proposition 6.5, it is easy to see that this is true if the word ‘immersed’ is replaced with ‘embedded’. In general, using Scott’s theorem on subgroup separability of surface groups [scott1, scott2], we can lift an immersed surface to an embedded one, and thus obtain an affirmative answer: